OpenStudy (anonymous):

How to solve this equation sin4x-2sinx=sinx :D

OpenStudy (mathmale):

\[\sin 4x~ or~ \sin ^{4}x~??\]Need some clarification first! Which of the following do you mean?

OpenStudy (mathmale):

You could, if you wanted, take a screen shot of the original problem and upload that image to OpenStudy.

OpenStudy (anonymous):

\[\sin 4x -\sin 2x=\sin x \] :D